OpenStudy (anonymous):

Which expression is the simplest form of sqrt m^20? a. m^2sqrtm^5 b. m^4sqrtm^5 c. m^4 d. m^10

OpenStudy (anonymous):

D. m^20 = (m^10)^2, so the sqrt(m^20) = m^10